A car travels at a speed of 48mph, another car leaves one hour later at a speed of 72mph. How long will it take for the second car to catch up to the first car?
Answer by mananth(16949)

speed of first car= 48 mph
speed of second car = 72 mph
second car leaves 1 hour late.
so when second car leaves first car has traveled 48 miles
catch up distance = 48 miles
same direction so catch up speed = 72-48 = 24 mph
t=d/r
t=48/24
t= 2 hours
catch up time = 2hours